Two of the top grossing concert tours were by a jazz band and a rock band. Together the two tours visited 174 cities. The jazz band visited 96 cities more than the rock band.
How many cities did each group visit.
Jazz band _____
Rock band _____

Answer by mastermath(14)

Let the number of cities visited by Rock band be x.
The jazz band visited 96 cities more than the rock band.
So, the number of cities visited by Jazz band will be (x+96).
Together the two tours visited 174 cities.
So, add them together.
Rock + Jazz = 174
x+(x+96)=174
2x+96=174
Now, subtract 96 on both the sides
2x+96-96=174-96
2x=78
Now, divide with 2 on both the sides
x=39
Thus, x+96 = 39+96 =135.
So, the number of cities visited by Rock band is 39.
And the number of cities visited by Jazz band is 135.